OpenStudy (anonymous):

Multiply by simplifying and factoring. Assume all expressions under radicals represent nonnegative numbers. ^3sqrt y^4 ^3sqrt16y^5=

OpenStudy (anonymous):

\[\sqrt[3]{y^4} \sqrt[3]{16y^5} =\]

OpenStudy (anonymous):

y^(4/3) * 16^(1/3) * y^(5/3) = cuberoot 16 * y^3 i thin

OpenStudy (anonymous):

that' s an awkward one!

OpenStudy (anonymous):

one way to check if this is right is to let y be say 4 and calculate values of the two expressions to see if they are equal - ill try that now

OpenStudy (anonymous):

yep - with x=4 they both work out to 161.2699 so cuberoot 16 y^3 is right